DISCLAMER AT THE END OF THIS FILE....

July 1997,

Well, about one year ago, I desided to release the sources of my attempt to make an IFF ILBM Converter, see text below. So, I released it on a BBS of a friend of mine. Wondering through old suorces, I came across this piece of programming. Realizing that there’s a lot of new addition to the AmiNet on a weekly basis, I thought it would be a good idea to release the sources on AmiNet as well. Even though it’s a year later. The Amiga community has not parished :). This is my little gift to all of you.

September 1996,

Historical overview This package is a give away. Use it the way you see fit. I started this program sometime in 1993, I think. Anyway, I started to code the whole lot in assembler. This turned out to be a pain. You don’t need great speed to use a GUI. Most of the time the program is idle, waiting for you to do something. So, after a long time I diceded that C might be a good alternative. One problem though, I didn’t know how to program in the beautiful language of C. So, the challenge was to teach myself the beautiful language, C. The program grew and grew and got just too big to be in one module. So a redesign was in order. This was more or less the thrid attempt to write an IFF ILBM Converter. By now, Commodore got into trouble :( and the will to continue was laking. So, little work was done in last 1 to 1.5 years. Now that I see absolutly no hope that the Amiga will ever be anything I decided to give away all my sources for free. My first intention was to make the finished version shareware, but since I will not finish the program and it is just no good sitting on my hard disk, I thought by giving it away I might do some good.

The sources are developed with SASC 6.50 C Compiler from SAS Institute,

Inc. and with Devpac Macro Assembler 3.04 from HiSoft.

ONE BIG NOTE!!! The save function can be dangerous to use. It has *not* been tested. The save copper option should work fine, but the save structure has been changed and I do not know what the consequences are for saving the copper lists. BE WARNED!!!

So, take a look at the source, learn from it or use the code. It is yours to use. If you like it and use it in anyway, please send me a postcard or an e-mail. This way I know someone made good use of it.

address:  e-mail (valid until around June/Jul 1998)

      Gerben Venekamp  venekamp@wins.uva.nl

Nuyenburglaan 9

1701 AJ Heerhugowaard

The Netherlands

Included files:

AslManagement.c  2728 ----rw-d 07-Jan-97 22:12:36
AslManagement.c.info  1573 ----rw-d 07-Jan-97 22:12:36
Build.info  1740 ----rw-d 07-Jan-97 22:12:36
Debug  2632 ----rwed 07-Jan-97 22:12:37
Debug.info  1347 ----rw-d 07-Jan-97 22:12:37
DecomPic.a  2232 ----rwed 07-Jan-97 22:12:37
DecomPic.a.bak  2340 ----rwed 07-Jan-97 22:12:37
DecomPic.a.info  1778 ----rw-d 07-Jan-97 22:12:37
DiskFonts.c  1142 ----rw-d 07-Jan-97 22:12:37
DiskFonts.c.info  1573 ----rw-d 07-Jan-97 22:12:37
Edit.info  1737 ----rw-d 07-Jan-97 22:12:37
ErrorText.h  1898 ----rwed 07-Jan-97 22:12:37
ErrorText.h.info  1613 ----rw-d 07-Jan-97 22:12:38
Find.info  1741 ----rw-d 07-Jan-97 22:12:38
Gadgets.c  2909 ----rw-d 07-Jan-97 22:12:38
Gadgets.c.info  1573 ----rw-d 07-Jan-97 22:12:38
Gadgets.h  6084 ----rw-d 07-Jan-97 22:12:38
Gadgets.h.info  1613 ----rw-d 07-Jan-97 22:12:38
Gadgets enum.h  642 ----rw-d 07-Jan-97 22:12:38
Gadgets enum.h.info  1613 ----rw-d 07-Jan-97 22:12:38
Gadgets enum.i  209 ----rwed 07-Jan-97 22:12:38
Gadgets enum.i.bak  160 ----rwed 07-Jan-97 22:12:38
Gadgets enum.i.info  1563 ----rw-d 07-Jan-97 22:12:38
GenAm.opts  66 ----rwed 07-Jan-97 22:12:39
GenAm.opts.info  1623 ----rw-d 07-Jan-97 22:12:39
HandleIntuiMessages.c  18531 ----rwed 07-Jan-97 22:12:39
HandleIntuiMessages.c.info  1573 ----rw-d 07-Jan-97 22:12:39
IFFConverter 23940 ----rwed Today 18:44:25
IFFConverter.c  14888 ----rw-d 07-Jan-97 22:12:39
IFFConverter.c.info  1573 ----rw-d 07-Jan-97 22:12:39
IFFConverter.h  5897 ----rw-d 07-Jan-97 22:12:39
IFFConverter.h.info  1613 ----rw-d 07-Jan-97 22:12:39
IFFConverter.info  1548 ----rw-d 07-Jan-97 22:12:40
IFFConverter.lnk  121 ----rwed 07-Jan-97 22:12:40
IFFStartup.a  2766 ----rwed 07-Jan-97 22:12:40
IFFStartup.a.bak  2766 ----rwed 07-Jan-97 22:12:40
IFFStartup.a.info  1778 ----rw-d 07-Jan-97 22:12:40
LoadSavePicture.c  22988 ----rwed 07-Jan-97 22:12:40
LoadSavePicture.c.bak  13774 ----rwed 07-Jan-97 22:12:41
LoadSavePicture.c.info  1573 ----rw-d 07-Jan-97 22:12:41
Manual.info  1922 ----rw-d 07-Jan-97 22:12:41
MemoryManagement.c  4436 ----rw-d 07-Jan-97 22:12:41
MemoryManagement.c.info  1573 ----rw-d 07-Jan-97 22:12:41
Misc.c  4384 ----rw-d 07-Jan-97 22:12:41
Misc.c.info  1573 ----rw-d 07-Jan-97 22:12:41
OpenCloseLibs.c  3906 ----rwed 07-Jan-97 22:12:41
OpenCloseLibs.c.info  1573 ----rw-d 07-Jan-97 22:12:42
PaletteMaintenance.c  5335 ----rw-d 07-Jan-97 22:12:42
PaletteMaintenance.c.info  1573 ----rw-d 07-Jan-97 22:12:42
README  7679 ----rwed Today  19:13:10
README.info  1154 ----rw-d Today  19:07:30

continued on next page contunued from previous page

README.dvi  10940 ----rwed Today  21:07:13
README.ps  158624 ----rwed Today  21:11:30
README.tex  10135 ----rwed Today  21:14:28
SavePicture.a  17463 ----rwed 07-Jan-97 22:12:42
SavePicture.a.bak  17459 ----rwed 07-Jan-97 22:12:43
SavePicture.a.info  1788 ----rw-d 07-Jan-97 22:12:43
SCoptions 164 ----rwed Today 18:41:55
SCoptions.info  1621 ----rw-d 07-Jan-97 22:12:43
smakefile  3043 ----rw-d 07-Jan-97 22:12:43
smakefile.info  1617 ----rw-d 07-Jan-97 22:12:43
WindowScreenMaintenace.c  20661 ----rw-d 07-Jan-97 22:12:43
WindowScreenMaintenace.c.info  1573 ----rw-d 07-Jan-97 22:12:43
Manual  Dir ----rwed 07-Jan-97 22:12:44

65 files - 1 directory - 1009 blocks used

Directory ”Manual” on Wednesday 16-Jul-97

Manual.guide  30104 ----rwed 07-Jan-97 22:12:44
Manual.guide.info  1796 ----rw-d 07-Jan-97 22:12:44

2 files - 65 blocks used

TOTAL: 67 files - 1 directory - 1074 blocks used

DISCLAMER:

This software pakage is provided AS IS and is free of any charge. No liability is taken. Any damage by this software or use of this software is yours. No problems have occoured on my system. This software has been developed with KS3.0 or higher in mind.